# Pulling apart the mechanisms that lead to jammed knitted fabrics

### v1

@@ IV Reducing or Inducing Jamming
To understand how changing yarn properties and manufacturing parameters affects jamming, we must first quantify fabric jamming. We identify two different metrics for measuring the strength of jamming that can be measured from any stress-strain curve. First is the slope of the jammed part of the curve at low strain, which we estimate with the secant modulus:
@@
Next, we examine conditions under which fabrics become jammed. In analogy with the jamming phase diagram for athermal granular systems [33], we propose that fabric jamming occurs for sufficiently high stitch density (which occurs when the amount of yarn forming the stitch is small relative to the yarn radius) and low enough applied stress, as shown in Fig. 1(c). Notably, the length of yarn per stitch $L$ and yarn radius $r$ are not independent and both vary with the “stitch size” setting on industrial knitting machines, as described in Supplementary Section V.
The secant modulus $Y_{\rm sec}$ and offset stress $\sigma_{0}$ were calculated for all experimental fabric samples in both directions, shown in Fig. 2c,d. Generally, increasing $L/r$ (“stitch size”) leads to less jammed fabrics as seen by decreasing secant modulus $Y_{\rm sec}$ and offset stress $\sigma_{0}$. This is consistent with qualitative examination of the stress-strain curves in Fig. S1. Unfortunately, yarn inhomogeneities lead to large error in $L/r$ and makes determining an exact relationship between the jamming metrics and fabric parameters difficult. Fits for this data, and all other fits contained in Fig. 2, can be found in Table S1. Within this error, however, we can see that the fabrics are more jammed and have larger $Y_{\rm sec}$ in the $y$-direction than the $x$-direction, as seen in Fig. 2c.
@@
Changing the yarn radius alone has a much smaller affect on the two jamming metrics (Fig. 2e,f). Although increasing the yarn radius leads to increasingly-jammed fabrics, as seen by an increase in both $Y_{\rm sec}$ and $\sigma_{0}$, the response is distinct from changes to $L$, suggesting that additional parameters are needed to characterize the jamming diagram. Since the exact way the core radius changes with machine tension is unknown, we also varied the yarn radius with a fixed ratio of $r_{\rm core}/r=0.45$. The ratio $r_{\rm core}/r$ seems to have a larger effect on the jamming behavior than changing the yarn radius $r$ alone (seen in Fig. 2e,f and Fig. S4) and the stiff jammed regime is larger as measured by both the secant modulus and offset stress. We can rationalize the augmented jamming response to changes in $r_{\rm core}$ in relation to the other yarn lengthscales given that the compression stiffness has a nonlinear dependence on $r_{\rm core}$. It is possible that the fixed ratio of core radius to yarn radius may induce more jamming with a larger secant modulus $Y_{\rm sec}$ than the length of yarn per stitch, but these simulations for large $r$ (small $L/r$) became numerically unstable.
Figure 3: Energy gap for the contact (pink) and bending (blue) energies as a function of length of yarn per stitch for compression scaling constant $k=0.1$ mN/mm2 pulled in the x-direction (a) and the y-direction (b). The light green region shows where in length of yarn per stitch parameter space the fabric moves from jammed ($\Delta E_{c}>5*10^{-5}$ J) to un-jammed ($\Delta E_{c}<5*10^{-5}$ J). Both log-linear and rest-energy rescaling depictions of these plots are given in Fig. S15.
@@ V Energy Markers for Jamming
Varying the length of yarn per stitch $L$ for different values of the compression scaling constant $k$ indicates the presence of three regimes within the energy difference: jamming, a transition region, and un-jammed fabric. For extension in both the $x$- and $y$-directions, we see all three regimes, though they are present in different quantities: jamming in the $x$-direction has an energy gap in the contact energy whereas jamming in the $y$-direction has an energy gap in the bending energy (Fig. 3). The linear regime of the curve in Fig. 4a indicates jamming, when the energy-minimizing configuration of the fabric does not correspond to the minimum contact energy (for the $x$-direction) or the minimum bending energy (for the $y$-direction). The horizontal region along the $x$-axis indicates an unjammed or nearly unjammed state, when the energy-minimizing configuration also corresponds the minimum contact energy (for the $x$-direction) or the minimum bending energy (for the $y$-direction). We describe the onset of unjamming, the first stitch length with an energy gap below our error threshold, as $(L/r)_{\rm trans}$ (Summplementary Section VII). We characterize the distance between the $x$-intercept of the linear jammed regime and $(L/r)_{\rm trans}$ as $\Delta L/r$. $\Delta L/r$ shows that the transition region from jammed to unjammed behavior is small for low compression constants and becomes larger for larger compression constants, shown in Fig. 4b.
The exact jamming transition lies within this region. Small compression constants, $k\lesssim 0.1mN/mm^{2}$, are near one of the incompressible limits, $k\to 0$, where the compression force in the outer core of the core-shell model becomes zero and the inner core remains incompressible. At large $k$, it becomes difficult to determine if fabrics near the jamming transition region, as defined by the energy gap, are jammed or un-jammed from the stress-strain plots.
Figure 4: (a) Contact energy gap as a function of length of yarn per stitch for compression scaling constant $k=2$ $mN/mm^{2}$ for fabrics stretched in the x-direction. The onset of un-jamming is marked with the vertical dashed line labelled $(L/r)_{trans}$. $\Delta L/r$ is a measure of the size of the jamming transition and defined as the distance between the x-intercept of the linear fit to the jammed regime (the dotted line) and $(L/r)_{trans}$. The linear fit is described in Table S4. Error in the energy calculation is described in Supplementary Section VII and Table S2. (b) The jamming transition size as a function of the compression scaling constant $k$. The jamming transition grows larger as $k$ increases and the yarn moves further from the incompressible limit. Error in $\Delta L/r$ is described in Fig. 3. (c) Phase diagram showing the jammed and unjammed regimes for different lengths of yarn per stitch $L$, compression scaling constants $k$, and bending moduli $B$. The top center plot is a phase space diagram showing $(L/r)_{trans}$ as a function of the compression scaling constant $k$ and the bending modulus $B$ for the x-direction (orange) and the y-direction (green). The lower left shows a cut of this 3D phase space with varying $k$ and the lower right is a cut of varying $B$. The pink regimes are jammed in both the x- and y-directions, the orange regimes are jammed only in the y-direction, and the green regimes are unjammed. The onset of unjamming, $(L/r)_{trans}$, is given by the solid lines, where an orange solid line is the end of jamming in the x-direction and a green solid line is the end of jamming in the y-direction. Increasing $k$ leads to a wider gap of $(L/r)_{trans}$ between the x- and y-directions whereas increasing $B$ closes that gap.
@@
Figure 5: (a) Stitch cell ratio $a_{y}/a_{x}$ as a function of $(L/r)^{2}$ for two simulation runs ($k=0.1$ $N/mm^{2}$ and $k=1.0$ $N/mm^{2}$) in closed symbols and experimental data in open symbols. For the simulation data, there are three regimes: fabrics jammed in both extension directions (pink), fabrics jammed only in the y-direction (orange), and fabrics that are not jammed in either direction (green). The transition from jammed in both directions to jammed only in the y-direction gives the peak of the aspect ratio curve. The experimental data is not inconsistent with this trend; though the data is sparse and the error bars are large, this transition is expected between $375<(L/r)^{2}\leq 423$ in the experimental data from the stress-strain plots in Fig. S1. Due to manufacturing constraints, the experimental data does not display a sample that is un-jammed in both extension directions. (b) Simulation values for the area strain as a function of strain in the y-direction for fabrics of different lengths of yarn per stitch $L$ stretched in the y-direction. Unjammed fabrics, $L>11.6$ mm, have non-monotonic area strain that dips below zero, indicating that the zero-force configuration is not the minimum area configuration when they are stretched in the y-direction. Jammed fabrics, by contrast, have a zero-force configuration that is also the minimum area configuration. Area strain for the x-direction is given in Fig. S13, experimental area strain is given in Fig. S14, and these results are discussed in more detail in Supplementary Section VI. (e) Photograph of the experimental sample made at the smallest sampled machine tension setting (10.5). (f) Photograph of the experimental sample made at the largest sampled machine tension setting (13). These photographs show how changing the machine tension setting leads to visible differences in the length of yarn per stitch, the area of the stitches, the stitch aspect ratio, and the yarn radius. These effects are quantified in Supplementary Section II. Photo credit: A. Patrick Cachine.
@@
For jammed fabrics stretched in the $y$-direction, $\Delta E_{c}$ is zero and the contact energy is at a minimum even though the total contact energy is smaller than the total bending energy (Fig. 3b), implying that yarn contacts dominate this regime [15]. For increasing compression scaling constant $k$, the onset of unjamming, $(L/r)_{trans}$, moves to larger and larger values of $L$, shown in Fig. 4c. For small values of $k$ (the incompressible regime), the transition coincides for both the $x$- and $y$-directions. As $k$ increases the $y$-direction becomes “more jammed” than the $x$-direction; an increasing range of lengths of yarn per stitch values becomes jammed in the $y$-direction but not jammed in the $x$-direction.
@@
As the bending modulus increases, the onset of unjamming, $(L/r)_{trans}$, remains nearly constant for fabrics stretched in the $x$-direction. There is a slight decrease in $(L/r)_{trans}$ for relatively large compression constant values ($k>1$ $mN/mm^{2}$), seen in Fig. 4 and Fig. S6. By comparison, the bending modulus $B$ has a minor effect on the jamming behavior, with larger values of $B$ corresponding to jamming occurring at smaller stitch lengths $L/r$ (i.e. denser fabrics); this decrease in $(L/r)_{\rm trans}$ is larger for $x$-direction jamming. As the bending stiffness of the yarn increases, the compression stiffness plays a reduced role in determining yarn geometry. Consequently, $(L/r)_{\rm trans}$ for $x$- and $y$-direction converge to similar values for increasing $B$, similar to the $k\to 0$ limit. As shown in Fig. S6, $(L/r)_{trans}$ in the $x$-direction never becomes significantly larger than $(L/r)_{trans}$ in the $y$-direction, implying that fabrics that are jammed in the $x$-direction and not the $y$-direction are very unlikely to exist or impossible to manufacture.
Figure 6: (a-e) Stitch contact volumes shown for a jammed fabric with a length of yarn per stitch $L=10.5$ mm. (a) The self-contact shown in green. (b) The nearest-neighbor contact with stitches in the same row but different columns shown in magenta. (c) The nearest-neighbor contact with stitches in the same column a row above (darker blue) or below (cyan). (d) The next-nearest-neighbor contact with stitches in the same column two rows above or below in yellow. (e) All contact volumes shown together in the stitch for a jammed fabric with length of yarn per stitch $L=10.5$ mm. (f) Contact volumes shown for an un-jammed fabric with length of yarn per stitch $L=12.2$ mm. The contact volumes for (a-f) look fuzzy at the edges because they are made of a semi-transparent clouds of points. (g) Strain in the z-direction as a function of strain in the x-direction for an un-jammed ($L=12.2$ mm in filled circles) and jammed ($L=10.5$ mm in open circles) fabric being stretched in the x-direction. The grey horizontal line is zero-strain. (h) Magnitude of the ratio of total contact forces in the x- and y-directions for zero-strain configurations of fabrics with different lengths of yarn per stitch $L$. Near the jamming transition, we see that the contact force in the y-direction becomes larger than the contact force in the x-direction. The dashed blue horizontal line represents equal contact forces in the x- and y-directions. Error in contact force is described in Table S3.
@@ VI Anisotropic Markers in Fabric Dimensions
Another anisotropic marker of jamming arises in the shape of the stitch unit cell. For fabrics jammed in the $y$-direction, the zero-strain configuration is the minimum stitch cell area configuration over the entire strain sweep (Fig. 5b). For un-jammed fabrics, the zero-strain configuration is not the minimum stitch cell area configuration; as the fabric is pulled in the $y$-direction, the stitch cell area decreases and then increases, shown in Fig. 5. This indicates that fabrics jammed in the $y$-direction are geometrically confined and at the smallest possible stitch area for their set of yarn and manufacturing parameters. Fabrics stretched in the $x$-direction have a non-monotonic stitch cell area as a function of strain regardless of whether or not the fabric is jammed in that direction and simulated fabrics all take on the minimum stitch cell area configuration at zero-strain (Fig. S13). We chose to analyse the stitch area instead of the stitch volume because existing experimental procedure can determine area strain. However, experimental area strain results, calculated from the motion of internal pins as seen in Fig. 2e,f, have very large errors that obscure possible trends in area strain due to jamming. These results are in Fig. S14 and discussed in more detail in Supplementary Section VI.
The stitch cell aspect ratio of the zero-strain configuration also gives information on jamming. For small compression scaling constant $k$ in the in-compressible limit, the stitch cell aspect ratio $a_{y}/a_{x}$ as a function of $(L/r)^{2}$ reaches a maximum near the transition from jammed to un-jammed (Fig. 5a). For larger $k$, with softer transitions and a mismatch between jamming transitions in the $x$- and $y$-directions, the aspect ratio reaches a maximum when the $x$-direction becomes un-jammed. This is also true for other simulation results with mis-matched jamming transitions in the $x$- and $y$-directions. Experimental measurements of the fabric aspect ratio show a similar trend (Fig. 5a). Due to the small number of experimental samples (six) and the large variance in the measured yarn radius, it is difficult to determine if the maximum aspect ratio aligns exactly with the onset of un-jamming in the $x$-direction.

### v2

@@ IV Reducing or Inducing Jamming
To understand how changing yarn properties and manufacturing parameters affect jamming, we must first quantify fabric jamming. We identify two different metrics for quantifying the behavior around the jamming transition that can be measured from any stress-strain curve. First is the slope of the jammed part of the curve at low strain, which we estimate with the secant modulus:
@@
Next, we examine conditions under which fabrics become jammed. In analogy with the jamming phase diagram for athermal granular systems [33], we propose that fabric jamming occurs for sufficiently high stitch density (which occurs when the amount of yarn forming the stitch is small relative to the yarn radius) and low enough applied stress, as shown in Fig. 1c. Notably, the length of yarn per stitch $L$ and yarn radius $r$ are not independent and both vary with the “stitch size” setting on the M1 Plus software used to program our industrial knitting machine as described in Supplementary Section V.
The secant modulus $Y_{\rm sec}$ and offset stress $\sigma_{0}$ were calculated for all experimental fabric samples in both directions, shown in Fig. 2c,d. Generally, increasing $L/r$ (“stitch size”) leads to less jammed fabrics as seen by decreasing secant modulus $Y_{\rm sec}$ and offset stress $\sigma_{0}$. This is consistent with qualitative examination of the stress-strain curves in Fig. S1. We can see that the fabrics have larger $Y_{\rm sec}$ in the $y$-direction than the $x$-direction, implying that they are “more jammed” or harder to soften in the $y$-direction as seen in Fig. 2c. Fits for this data, and all other fits contained in Fig. 2, can be found in Table S1.
@@
Figure 3: Energy gap for the contact (coral) and bending (blue) energies as a function of length of yarn per stitch for compression scaling constant $k=0.1$ mN/mm2 pulled in the $x$-direction (a) and the $y$-direction (b). The light green region shows where in length of yarn per stitch parameter space the fabric moves from jammed ($\Delta E_{c}>5*10^{-5}$ J) to un-jammed ($\Delta E_{c}<5*10^{-5}$ J). Both log-linear and rest-energy rescaling depictions of these plots are given in Fig. S15.
Changing the yarn radius alone has a much smaller effect on the two jamming metrics (Fig. 2e,f). Although increasing the yarn radius leads to increasingly-jammed fabrics, as seen by an increase in both $Y_{\rm sec}$ and $\sigma_{0}$, the response is distinct from changes to $L$, suggesting that additional parameters are needed to characterize the jamming diagram. Since the exact way the core radius changes with “stitch size” is unknown, we also varied the yarn radius with a fixed ratio of $r_{\rm core}/r=0.45$. In simulations, the ratio $r_{\rm core}/r$ seems to have a larger effect on the jamming behavior than changing the yarn radius $r$ alone (seen in Fig. 2e,f and Fig. S4) and the stiff jammed regime is larger as measured by both the secant modulus and offset stress. We can rationalize the augmented jamming response to changes in $r_{\rm core}$ in relation to the other yarn lengthscales given that the compression stiffness has a nonlinear dependence on $r_{\rm core}$. It is possible that the fixed ratio of core radius to yarn radius may induce more jamming with a larger secant modulus $Y_{\rm sec}$ than the length of yarn per stitch, but these simulations for large $r$ (small $L/r$) became numerically unstable.
@@ V Energy Markers for Jamming
Changing the length of yarn per stitch $L$ for different values of the compression scaling constant $k$ indicates the presence of three regimes within the energy difference: jamming, a transition region, and un-jammed fabric. For extension in both the $x$- and $y$-directions, we see all three regimes, though the energy contributions differ for each direction. Jamming in the $x$-direction has an energy gap in the contact energy whereas jamming in the $y$-direction has an energy gap in the bending energy (Fig. 3). The linear regime of the curve in Fig. 4a indicates jamming when the energy-minimizing configuration of the fabric does not correspond to the minimum contact energy (for the $x$-direction) or the minimum bending energy (for the $y$-direction). The horizontal region along the $x$-axis indicates an unjammed or nearly unjammed state, when the energy-minimizing configuration also corresponds the minimum contact energy (for the $x$-direction) or the minimum bending energy (for the $y$-direction). We describe the onset of unjamming, the first stitch length with an energy gap below our error threshold, as $(L/r)_{\rm trans}$ (Supplementary Section VII). We characterize the distance between the $x$-intercept of the linear jammed regime and $(L/r)_{\rm trans}$ as $\Delta L/r$. $\Delta L/r$ shows that the transition region from jammed to unjammed behavior is small for low compression constants and becomes larger for larger compression constants, shown in Fig. 4b.
The exact jamming transition lies within this region. Small compression constants, $k\lesssim 0.1{\rm mN/mm^{2}}$, are near one of the incompressible limits, $k\to 0$, where the compression force in the outer core of the core-shell model becomes zero and the inner core remains incompressible. At large $k$, it becomes difficult to determine if fabrics near the jamming transition region, as defined by the energy gap, are jammed or un-jammed from the stress-strain plots.
@@
Figure 4: (a) Contact energy gap as a function of length of yarn per stitch for compression scaling constant $k=2~{}{\rm mN/mm^{2}}$ for fabrics stretched in the $x$-direction. The onset of un-jamming is marked with the vertical dashed line labelled $(L/r)_{\rm trans}$. $\Delta L/r$ is a measure of the size of the jamming transition and defined as the distance between the x-intercept of the linear fit to the jammed regime (the dotted line) and $(L/r)_{\rm trans}$. The linear fit is described in Table S4. Error in the energy calculation is described in Supplementary Section VII and Table S2. (b) The jamming transition region $\Delta L/r$ as a function of the compression scaling constant $k$. The jamming transition region becomes larger as $k$ increases and the yarn moves further from the incompressible limit, $k\rightarrow 0$. Error in $\Delta L/r$ is described in Fig. 3. (c) Phase diagram showing the jammed and unjammed regimes for different lengths of yarn per stitch $L$, compression scaling constants $k$, and bending moduli $B$. The top center plot is a phase space diagram showing $(L/r)_{\rm trans}$ as a function of the compression scaling constant $k$ and the bending modulus $B$ for the $x$-direction (orange) and the $y$-direction (green). The lower left shows a slice of this 3D phase space with varying $k$ and the lower right is a slice of varying $B$. The pink regimes are jammed in both the $x$- and $y$-directions, the orange regimes are jammed only in the $y$-direction, and the green regimes are unjammed. The onset of unjamming, $(L/r)_{\rm trans}$, is given by the solid lines, where an orange solid line is the end of jamming in the $x$-direction and a green solid line is the end of jamming in the $y$-direction. Increasing $k$ leads to a wider gap of $(L/r)_{\rm trans}$ between the $x$- and $y$-directions; whereas, increasing $B$ closes that gap.
@@
Figure 5: (a) Aspect ratio of a stitch $a_{y}/a_{x}$ as a function of $(L/r)^{2}$ for two simulation runs ($k=0.1~{}{\rm N/mm^{2}}$ and $k=1.0~{}{\rm N/mm^{2}}$) in closed symbols and experimental data in open symbols. For the simulation data, there are three regimes: fabrics jammed in both extension directions (pink), fabrics jammed only in the y-direction (orange), and fabrics that are not jammed in either direction (green). The transition from jammed in both directions to jammed only in the $y$-direction gives the peak of the aspect ratio curve. The experimental data is not inconsistent with this trend; though the data is sparse and the error bars are large, this transition is expected between $375<(L/r)^{2}\leq 423$ in the experimental data from the stress-strain plots in Fig. S1. Due to manufacturing constraints, the experimental data does not display a sample that is un-jammed in both extension directions. (b) Simulation values for the area strain as a function of strain in the y-direction for fabrics of different lengths of yarn per stitch $L$ stretched in the $y$-direction. Unjammed fabrics, $L>11.6~{}{\rm mm}$, have non-monotonic area strain that dips below zero, indicating that the zero-force configuration is not the minimum area configuration when they are stretched in the y-direction. Jammed fabrics, by contrast, have a zero-force configuration that is also the minimum area configuration. Area strain for the $x$-direction is given in Fig. S13, experimental area strain is given in Fig. S14, and these results are discussed in more detail in Supplementary Section VI. (e) Photograph of the experimental sample made at the smallest sampled machine tension setting (10.5). (f) Photograph of the experimental sample made at the largest sampled “stitch size” setting (13). These photographs show how changing the “stitch size” setting leads to visible differences in the length of yarn per stitch, the area of the stitches, the stitch aspect ratio, and the yarn radius. These effects are quantified in Supplementary Section II. Photo credit: A. Patrick Cachine.
@@
For jammed fabrics stretched in the $y$-direction, the change in the compression energy $\Delta E_{c}$ is zero, and we see that the contact energy is at a minimum (Fig. 3b) even though the total contact energy is smaller than the total bending energy [15]. This implies that yarn contacts dominate in this regime. For increasing compression scaling constant $k$, the onset of unjamming, $(L/r)_{\rm trans}$, moves to larger and larger values of $L$, shown in Fig. 4c. For small values of $k$ (the incompressible regime), the transition coincides for both the $x$- and $y$-directions. As $k$ increases, the $y$-direction becomes “more jammed” than the $x$-direction. We find that the range of $L$ values that are jammed in the $y$-direction but not jammed in the $x$-direction becomes larger as $k$ increases.
@@
As the bending modulus increases, the onset of unjamming, $(L/r)_{\rm trans}$, remains nearly constant for fabrics stretched in the $x$-direction. There is a slight decrease in $(L/r)_{\rm trans}$ for relatively large compression constant values ($k>1\,{\rm mN/mm^{2}}$), seen in Fig. 4 and Fig. S6. By comparison, the bending modulus $B$ has a minor effect on the jamming behavior, with larger values of $B$ corresponding to jamming occurring at smaller stitch lengths $L/r$ (i.e. denser fabrics). This decrease in $(L/r)_{\rm trans}$ is larger for $y$-direction jamming. As the bending stiffness of the yarn increases, the compression stiffness plays a reduced role in determining yarn geometry. Consequently, $(L/r)_{\rm trans}$ for $x$- and $y$-direction converge to similar values for increasing $B$, similar to the $k\to 0$ limit. As shown in Fig. S6, $(L/r)_{\rm trans}$ in the $x$-direction never becomes significantly larger than $(L/r)_{\rm trans}$ in the $y$-direction, implying that fabrics that are jammed in the $x$-direction and not the $y$-direction are very unlikely to exist or impossible to manufacture.
@@ VI Anisotropic Markers in Fabric Dimensions
Another anisotropic marker of jamming arises in the shape of the stitch unit cell. For fabrics jammed in the $y$-direction, the zero-strain configuration is the minimum area-per-stitch over the entire strain sweep (Fig. 5b). For un-jammed fabrics, the zero-strain configuration is not the minimum area-per-stitch; as the fabric is pulled in the $y$-direction, the area-per-stitch decreases and then increases, shown in Fig. 5. This indicates that fabrics jammed in the $y$-direction are geometrically confined and at the smallest possible area-per-stitch for their set of yarn and manufacturing parameters. Fabrics stretched in the $x$-direction have a non-monotonic area-per-stitch as a function of strain regardless of whether or not the fabric is jammed in that direction and simulated fabrics all take on the minimum area-per-stitch at zero-strain (Fig. S13). We chose to analyze the stitch area instead of the stitch volume because existing experimental procedure can determine area strain. However, experimental area strain results, calculated from the motion of internal pins as seen in Fig. 2e,f, have very large errors that obscure possible trends in area strain due to jamming. These results are in Fig. S14 and discussed in more detail in Supplementary Section VI.
The aspect ratio of the stitch cell for the zero-strain configuration also gives information on jamming. For small compression scaling constant $k$ in the in-compressible limit, the aspect ratio $a_{y}/a_{x}$ as a function of $(L/r)^{2}$ reaches a maximum near the transition from jammed to un-jammed (Fig. 5a). For larger $k$, with softer transitions and a mismatch between jamming transitions in the $x$- and $y$-directions, the aspect ratio reaches a maximum when the $x$-direction becomes un-jammed. This is also true for other simulation results with mis-matched jamming transitions in the $x$- and $y$-directions. Experimental measurements of the fabric aspect ratio show a similar trend (Fig. 5a). Due to the small number of experimental samples (six) and the large variance in the measured yarn radius, it is difficult to determine if the maximum aspect ratio aligns exactly with the onset of un-jamming in the $x$-direction.